1. 1:10 ::X:50. Determine the value of X.

2. While working in the pediatric clinic, the nurse weighs a 5-year-old child as 20.5 kg. The mother asks, “How much is that in pounds?” Provide the child’s weight to the nearest whole pound

3. The nurse is tallying intake for a client with congestive heart failure who is on a 1,000-mL fluid restriction. The client has had 1 cups of coffee, 1/2 cup of water, 2 oz juice, 30% of a 10 oz cup of soup, and half a pint of milk. How much additional fluid in milliliters can the client consume for the remainder of the day?
Answer #1

1.  

1/10-x=50

One solution was found :

                   x = -499/10 = -49.900

Rearrange:

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation :

                     1/10-x-(50)=0

Step by step solution :

Step  1  :

             1
 Simplify   ——
            10

Equation at the end of step  1  :

    1          
  (—— -  x) -  50  = 0 
   10          

Step  2  :

Rewriting the whole as an Equivalent Fraction :

2.1   Subtracting a whole from a fraction

Rewrite the whole as a fraction using  10  as the denominator :

         x     x • 10
    x =  —  =  ——————
         1       10  

Equivalent fraction : The fraction thus generated looks different but has the same value as the whole

Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator

Adding fractions that have a common denominator :

2.2       Adding up the two equivalent fractions
Add the two equivalent fractions which now have a common denominator

Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible:

 1 - (x • 10)     1 - 10x
 ————————————  =  ———————
      10            10   

Equation at the end of step  2  :

  (1 - 10x)    
  ————————— -  50  = 0 
     10        

Step  3  :

Rewriting the whole as an Equivalent Fraction :

3.1   Subtracting a whole from a fraction

Rewrite the whole as a fraction using  10  as the denominator :

          50     50 • 10
    50 =  ——  =  ———————
          1        10   

Adding fractions that have a common denominator :

3.2       Adding up the two equivalent fractions

 (1-10x) - (50 • 10)     -10x - 499
 ———————————————————  =  ——————————
         10                  10    

Step  4  :

Pulling out like terms :

4.1     Pull out like factors :

   -10x - 499  =   -1 • (10x + 499)

Equation at the end of step  4  :

  -10x - 499
  ——————————  = 0 
      10    

Step  5  :

When a fraction equals zero :

 5.1    When a fraction equals zero ...

Where a fraction equals zero, its numerator, the part which is above the fraction line, must equal zero.

Now,to get rid of the denominator, Tiger multiplys both sides of the equation by the denominator.

Here's how:

  -10x-499
  ———————— • 10 = 0 • 10
     10   

Now, on the left hand side, the  10  cancels out the denominator, while, on the right hand side, zero times anything is still zero.

The equation now takes the shape :
   -10x-499  = 0

Solving a Single Variable Equation :

5.2      Solve  :    -10x-499 = 0

Add  499 to both sides of the equation :
                      -10x = 499
Multiply both sides of the equation by (-1) :  10x = -499


Divide both sides of the equation by 10:
                     x = -499/10 = -49.900

One solution was found :

                   x = -499/10 = -49.900

2. 45.2 pounds  

Converting kilograms to pounds

There are 2.2046226218 pounds in 1 kilogram. So, if you want to do the calculation in your head, multiply your kilogram figure by 2.2 to get an approximate figure for pounds. Should you wish to convert from pounds to kilos, divide the figure by 2.2 for an approximation.

4. I cup coffee (standard cup ) = 250 ml

1/2 cup water (standard cup ) = 125 ml

2 OZ juice = 60ml

10 OZ cup of soup = 296 ml 30% of soup patient consumed which is equal to 88.8 ml

total consumption is = 250ml+125ml+60ml+88.8ml

=523.8ml

Remainder of the day = 1000ml - 523.8ml=476.2ml

client consume for the remainder of the day is 476.2ml